OpenStudy (anonymous):

An engineer orders 8,000 liters of solvent for a pollution control equipment at P1.50 per liter and for every one centavo subtracted from the cost per liter,500 more liters will be ordered,at what price per liter will the solvent dealer charge to give a maximum sale? a. 45 cents b. 83 cents c. 72 cents d. 67 cents

OpenStudy (anonymous):

we can write this as sales = (1.50 - x)(8000 + 500x) where x represents decrease by one centavo

OpenStudy (anonymous):

\[(1.50-x)(8000+500x)=-500 x^2-7250x+12000\] a parabola that opens down. your job is to find the vertex

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

im a filipino i dont know cents...

OpenStudy (anonymous):

w8 i try

OpenStudy (anonymous):

the units are unimportant. call them centavos is fine

OpenStudy (anonymous):

44.5 or 45 cents? when u say x represents decrease by 1 centavo

OpenStudy (phi):

We have to tweak sat's equation. A quick google search shows there are 100 centavos in a peso. So the equation should be \((150-x)(8000+500x)= \)price (centavos) This is the same as \[ -500x^2 +75000x-8000x+1200000= -500x^2 +67000x+1200000\] Now find the peak of the parabola using -b/2a where b= 67000 and a= -500 \[ \frac{-67000}{2\cdot -500} = 67 \text{ centavos }\] 67 is the discount. To answer "at what price per liter will the solvent dealer charge " we deduct 0.67 pesos from 1.50 pesos to get 0.83 pesos, or 83 centavos

OpenStudy (anonymous):

thanks, i didn't catch that. i made the erroneous assumption that "cent' was short for "centavo"

OpenStudy (phi):

I only noticed when I kept getting a negative x. Something is fishy here....